OpenStudy (anonymous):

By first differentiating (√(1-x))/(√(1+x)) obtain an expression dy/dx in terms of x. Hence show that the gradient of the normal to the curve at point (x,y) is (1+x)(√(1-x^2))

OpenStudy (anonymous):

\[\frac{ \sqrt{1-x} }{ \sqrt{1+x} }\]

ganeshie8 (ganeshie8):

gradient of normal = -1/(gradient of tangent) = \(-\dfrac{1}{\left(\frac{\sqrt{1-x}}{\sqrt{1+x}}\right)'}\)

OpenStudy (anonymous):

I found dy/dx i.e \[\frac{ -2 }{ (1+x^{2}) }\] How do I show that the gradient of the normal is equal to \[(1+x)\sqrt{1-x ^{2}}\]

ganeshie8 (ganeshie8):

dy/dx looks wrong. im(wolfram) getting : dy/dx = http://www.wolframalpha.com/input/?i=%28%5Cfrac%7B+%5Csqrt%7B1-x%7D+%7D%7B+%5Csqrt%7B1%2Bx%7D+%7D%29%27

ganeshie8 (ganeshie8):

taking the reciprocal of above and negating it gives the expression you looking for

OpenStudy (anonymous):

No, actually the question says just to derivate (1+x)/(1-x) and then use the answer to obtain the gradient of the normal.

ganeshie8 (ganeshie8):

yes, so ?

OpenStudy (anonymous):

The marking scheme I have says to use a chain rule. I don't know how to do that.

ganeshie8 (ganeshie8):

im not getting u friend :/

OpenStudy (anonymous):

Okay, I'll post the screenshot.

ganeshie8 (ganeshie8):

okay

ganeshie8 (ganeshie8):

how come the radical sign disappeared ?

ganeshie8 (ganeshie8):

doesnt that look like a clear typo ?

OpenStudy (anonymous):

I don't know. This is what the marking scheme says.

ganeshie8 (ganeshie8):

Gotcha :)

ganeshie8 (ganeshie8):

you're right, they want us to use chain rule lol

ganeshie8 (ganeshie8):

\(f(g(x)) = \sqrt{\dfrac{1-x}{1+x}}\)

ganeshie8 (ganeshie8):

\(g(x) = \dfrac{1-x}{1+x}\) \(f(x) = \sqrt{x}\)

ganeshie8 (ganeshie8):

\([f(g(x))]' = f'(g(x)) \times g'(x)\)

ganeshie8 (ganeshie8):

plug and chug ?

ganeshie8 (ganeshie8):

\(g(x) = \dfrac{1-x}{1+x} \implies g'(x) = \dfrac{-2}{(1+x)^2}\)

OpenStudy (anonymous):

Okay. I did that too. f'(x)= \[1/2\sqrt{x}\]

ganeshie8 (ganeshie8):

\(f(g(x)) = \sqrt{\dfrac{1-x}{1+x}} \) \(\implies [f(g(x))]' = \dfrac{1}{2\sqrt{\dfrac{1-x}{1+x}}} \times \left(\dfrac{-2}{(1+x)^2}\right) \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

I am still not getting that expression. All I got was: \[\sqrt{1-x} (1+x)^{3/2}\]

OpenStudy (anonymous):

^This was after taking the negative inverse of the gradient.

ganeshie8 (ganeshie8):

Lol you're done then

ganeshie8 (ganeshie8):

just rearrange that expression

ganeshie8 (ganeshie8):

\(\sqrt{1-x} (1+x)^{3/2} = \sqrt{1-x} (\sqrt{1+x})^{3} \\ =\sqrt{1-x}\sqrt{1-x} (\sqrt{1+x})^{2} = \sqrt{1-x^2}(1+x)\)

OpenStudy (anonymous):

O Lord, this is such a bad question. Too much simplification. How could have anyone done it during an exam with under 5 minutes per question? -_-

OpenStudy (anonymous):

@ganeshie8 : Thank you. Much appreciated! :)

ganeshie8 (ganeshie8):

hahah its more of algebra... not calculus lol

ganeshie8 (ganeshie8):

u wlc :)